The product of -2 and a number is the same as the total of 15 and the number find the number. How do i set it up and get the answer
Answer by Maths68(1474)

Let x is the number that we need to find
Product of -2 and x = (-2)(x) = -2x
Total of 15 and x = 15+x
Product of -2 and x = Total of 15 and x
(-2)(x)=x+15
-2x=x+15
-2x-x=15
-3x=15
-3x/3=15/3
-x=5
X=-5
Check
(-2)(x)=x+15
(-2)(-5)=-5+15
10=10